# Robust Score-Based Quickest Change Detection

### v3

@@ V Robust Quickest Change Detection in Unnormalized and Score-Based Models
$(Q_{\infty},Q_{1})=\operatorname*{arg\,min}_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{F}(R_{1}\|R_{\infty}).$ (14)
Thus, the distributions $Q_{\infty}\in\mathcal{G}_{\infty},Q_{1}\in\mathcal{G}_{1}$ are closest as measured through the pseudo distance of Fisher divergence (see (8)). We remark that since the observations are assumed to be high-dimensional, the densities are not precisely known. As a consequence, the stochastic boundedness condition of [17] or the KL divergence cannot be used here to define the least favorable pair. On the other hand, the Fisher divergence can be computed for unnormalized and score-based models. In the rest of the paper, we assume that we can always find the least favorable pair $(Q_{\infty},Q_{1})$. In Section VII, we provide several examples where this assumption is valid.
@@
$\min_{T}\;\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathcal{L}_{\texttt{WADD}}(T)\geq\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\min_{T}\;\mathcal{L}_{\texttt{WADD}}(T)\geq\frac{\log\gamma}{\inf_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{KL}(P_{1}\|P_{\infty})}(1+o(1)),$ (15)
@@
$\frac{\log\gamma}{\inf_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{KL}(P_{1}\|P_{\infty})}$
provides a lower bound on the asymptotic performance of any algorithm. This lower bound can be achieved by a CUSUM-type algorithm $T^{*}$ if $\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T^{*}]\geq\gamma$ and the drift of this algorithm after the change is $\inf_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{KL}(P_{1}\|P_{\infty})$. Since we cannot calculate the KL divergence for unnormalized and score-based models, we use the Fisher divergence as a proxy measure. We note also that these two divergences coincide for certain Gaussian families [15].
@@
$z(x)\overset{\Delta}{=}\mathcal{S}_{\texttt{H}}(x,Q_{\infty})-\mathcal{S}_{\texttt{H}}(x,Q_{1}),$ (16)
@@
$Z(0)=0,$
$Z(n)\overset{\Delta}{=}(Z(n-1)+z(X_{n}))^{+},\;\forall n\geq 1.$
@@ VI Consistency and Delay and False Alarm Analysis of the RSCUSUM Algorithm
For the least favorable distributions $Q_{\infty},Q_{1}$ of Definition V.1, $\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})<\mathbb{D}_{F}(P_{\infty}\|Q_{1})$.
@@
If $\mathcal{G}_{\infty}$ contains only one element ($\mathcal{G}_{\infty}=\{P_{\infty}\}$), then clearly the least favorable pre-change distribution $Q_{\infty}$ is equal to $P_{\infty}$ and $\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})=0$. Then, Assumption VI.1 follows from the disjointness of $\mathcal{G}_{\infty},\mathcal{G}_{1}$ given in Assumption IV.3 and the regularity conditions given in Assumption IV.4.
@@
$\mathcal{G}_{\infty}\overset{\Delta}{=}\biggl{\{}x\mapsto\sum_{i=1}^{m}\alpha_{i}p_{i}(x):\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0\biggr{\}}.$ (18)
If $\mathbb{D}_{F}(P_{i}\|Q_{\infty})<\mathbb{D}_{F}(P_{i}\|Q_{1})$ for all $1\leq i\leq m$, then Assumption VI.1 holds for any $P_{\infty}\in\mathcal{G}_{\infty}$.
@@
We use $C_{R}$ to denote the term $\mathbb{E}_{X\sim R}\left[\frac{1}{2}\left\|\nabla_{x}\log r(X)\right\|_{2}^{2}\right]$ for any distribution $R$. Then
$\begin{split}\mathbb{E}_{X\sim R}[z(X)]=\mathbb{E}_{X\sim R}&[\mathcal{S}_{\texttt{H}}(X,Q_{\infty})-\mathcal{S}_{\texttt{H}}(X,Q_{1})]=\mathbb{D}_{\texttt{F}}(R\|Q_{\infty})-C_{R}-\mathbb{D}_{\texttt{F}}(R\|Q_{1})+C_{R}=\mathbb{D}_{\texttt{F}}(R\|Q_{\infty})-\mathbb{D}_{\texttt{F}}(R\|Q_{1}),\end{split}$
@@
$=\sum_{i=1}^{m}\alpha_{i}\mathbb{E}_{P_{i}}[z(X)]=\sum_{i=1}^{m}\alpha_{i}\bigg{(}\mathbb{D}_{F}(P_{i}\|Q_{\infty})-\mathbb{D}_{F}(P_{i}\|Q_{1})\bigg{)}$ (19)
It is given that $\mathbb{D}_{F}(P_{i}\|Q_{\infty})-\mathbb{D}_{F}(P_{i}\|Q_{1})<0$ for each $1\leq i\leq m$ and that $\alpha_{i}\geq 0$ for all $1\leq i\leq m$. Thus, $\mathbb{D}_{F}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1})$ can be written as the sum of $m$ nonpositive terms, at least one of which is strictly negative. ∎
@@ VI-A Reverse Triangle Inequality for Fisher Divergence
$=\int\frac{1}{2}\big{\|}\nabla\log\bigl{(}\xi q_{1}+(1-\xi)r_{1}\bigr{)}-\nabla\log q_{\infty}\big{\|}^{2}$
$\qquad\qquad\qquad\qquad\qquad\bigl{(}\xi q_{1}+(1-\xi)r_{1}\bigr{)}dx.$
Clearly $\mathcal{L}(\xi)$ is minimized at $\xi=1$, and $\frac{\partial\mathcal{L}(\xi)}{\partial\xi}\mid_{\xi=1^{-}}\leq 0$. Let $\mathcal{L}^{\prime}(\xi)=\frac{\partial\mathcal{L}(\xi)}{\partial\xi}$, we have
$\mathcal{L}^{\prime}(\xi)=\int\frac{1}{2}\bigl{(}q_{1}-r_{1}\bigr{)}\big{\|}\nabla\log q_{\xi}-\nabla\log q_{\infty}\big{\|}^{2}dx+\int q_{\xi}\nabla\left(\frac{q_{1}-r_{1}}{q_{\xi}}\right)^{T}\bigl{(}\nabla\log q_{\xi}-\nabla\log q_{\infty}\bigr{)}dx.$
@@
$\frac{1}{2}r_{1}\big{\|}\nabla\log q_{1}-\nabla\log q_{\infty}\big{\|}^{2}$
$=\frac{1}{2}r_{1}\big{\|}\nabla\log q_{1}-\nabla\log r_{1}\big{\|}^{2}+\frac{1}{2}r_{1}\big{\|}\nabla\log r_{1}-\nabla\log q_{\infty}\big{\|}^{2}+\underbrace{r_{1}\bigl{(}\nabla\log q_{1}-\nabla\log r_{1}\bigr{)}^{T}\bigl{(}\nabla\log r_{1}-\nabla\log q_{\infty}\bigr{)}}_{\text{term 1(a)}}.$ (21)
@@
$q_{1}\nabla\left(\frac{q_{1}-r_{1}}{q_{1}}\right)^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}=r_{1}\bigl{(}\nabla\log q_{1}-\nabla\log r_{1}\bigr{)}^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}.$ (24)
@@
The results follows since $\frac{\partial\mathcal{L}(\xi)}{\partial\xi}\mid_{\xi=1^{-}}\leq 0$. ∎
@@ VI-B Consistency of the RSCUSUM Algorithm
$\mathbb{E}_{\infty}\left[z(X_{1})\right]$ $=\mathbb{E}_{\infty}\left[\mathcal{S}_{\texttt{H}}(X_{1},Q_{\infty})-\mathcal{S}_{\texttt{H}}(X_{1},Q_{1})\right]=\mathbb{D}_{\texttt{F}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{F}(P_{\infty}\|Q_{1})<0,\;\text{and}$
@@
As in Theorem VI.3, we use $C_{R}$ to denote the term $\mathbb{E}_{X\sim R}\left[\frac{1}{2}\left\|\nabla_{x}\log r(X)\right\|_{2}^{2}\right]$ for any distribution $R$. Then
$\begin{split}\mathbb{E}_{\infty}&[\mathcal{S}_{\texttt{H}}(X_{1},Q_{\infty})-\mathcal{S}_{\texttt{H}}(X_{1},Q_{1})]=\mathbb{D}_{\texttt{F}}(P_{\infty}\|Q_{\infty})-C_{P_{\infty}}-\mathbb{D}_{\texttt{F}}(P_{\infty}\|Q_{1})+C_{P_{\infty}}=\mathbb{D}_{\texttt{F}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{F}}(P_{\infty}\|Q_{1}),\end{split}$
@@
$\begin{split}\mathbb{E}_{1}&[\mathcal{S}_{\texttt{H}}(X_{1},Q_{\infty})-\mathcal{S}_{\texttt{H}}(X_{1},Q_{1})]=\mathbb{D}_{\texttt{F}}(P_{1}\|Q_{\infty})-C_{P_{1}}-\mathbb{D}_{\texttt{F}}(P_{1}\|Q_{1})+C_{P_{1}}\geq\mathbb{D}_{\texttt{F}}(Q_{1}\|Q_{\infty}),\end{split}$
@@ VI-C False Alarm Analysis of the RSCUSUM Algorithm
$z_{\rho}(x)\overset{\Delta}{=}\rho z(x)=\rho\bigr{(}\mathcal{S}_{\texttt{H}}(x,Q_{\infty})-\mathcal{S}_{\texttt{H}}(x,Q_{1})\bigr{)}.$ (27)
@@
for some $\rho$ which satisfies the inequality of Equation 26. To satisfy the constraint of $\mathbb{E}_{\infty}[T_{\texttt{RSCUSUM}}]\geq\gamma$, it is enough to set the threshold $\omega=\frac{\log\gamma}{\rho}$.
@@
As a result, by setting $\omega=\frac{\log\gamma}{\rho^{*}}$, we get
@@ VI-D Delay Analysis of the RSCUSUM Algorithm
$\mathcal{L}_{\texttt{WADD}}$ $(T_{\texttt{RSCUSUM}})\sim\mathcal{L}_{\texttt{CADD}}(T_{\texttt{RSCUSUM}})\sim\mathbb{E}_{1}[T_{\texttt{RSCUSUM}}]\sim\frac{\omega}{\mathbb{D}_{\texttt{F}}(P_{1}\|Q_{\infty})-\mathbb{D}_{\texttt{F}}(P_{1}\|Q_{1})}\lesssim\frac{\omega}{\mathbb{D}_{\texttt{F}}(Q_{1}\|Q_{\infty})},\quad\text{as $\omega\to\infty$. }$
Furthermore, if we let $\omega=\frac{\log\gamma}{\rho}$ as in Theorem VI.7, then we have $\mathbb{E}_{\infty}[T_{\texttt{RSCUSUM}}]\geq\gamma$, and
$\mathcal{L}_{\texttt{WADD}}$ $(T_{\texttt{RSCUSUM}})\sim\mathcal{L}_{\texttt{CADD}}(T_{\texttt{RSCUSUM}})\lesssim\frac{\log\gamma}{\rho\mathbb{D}_{F}(Q_{1}\|Q_{\infty})},\quad\text{as $\gamma\to\infty$.}$
@@
In the above theorem, we have used the notation $g(c)\lesssim h(c)$ as $c\to c_{0}$ to indicate that $\lim\sup\frac{g(c)}{h(c)}\leq 1$ as $c\to c_{0}$ for any two functions $c\mapsto g(c)$ and $c\mapsto h(c)$.
Theorem VI.8 implies that the expected detection delay (EDD) increases at most linearly as the stop threshold $\omega$ increases for large values of $\omega$. Again, since the bound is valid for all those $\rho$ values for which $\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]\leq 1$, the bound is valid for the biggest such number, which is $\rho^{*}$. This gives
$\mathcal{L}_{\texttt{WADD}}(T_{\texttt{RSCUSUM}})\sim\mathcal{L}_{\texttt{CADD}}(T_{\texttt{RSCUSUM}})\lesssim\frac{\log\gamma}{\rho^{*}\mathbb{D}_{F}(Q_{1}\|Q_{\infty})},\quad\text{as $\gamma\to\infty$.}$ (30)

### v4

@@ V Robust Quickest Change Detection in Unnormalized and Score-Based Models
$(Q_{\infty},Q_{1})=\operatorname*{arg\,min}_{(R_{\infty},R_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{\texttt{{F}}}(R_{1}\|R_{\infty}).$ (14)
Thus, the distributions $Q_{\infty}\in\mathcal{G}_{\infty},Q_{1}\in\mathcal{G}_{1}$ are closest as measured through the pseudo distance of Fisher divergence (see (8)). We remark that since the observations are assumed to be high-dimensional, the densities are not precisely known. As a consequence, the stochastic boundedness condition of [17] or the KL divergence cannot be used here to define the least favorable pair. On the other hand, the Fisher divergence can be computed for unnormalized and score-based models. In the rest of the paper, we assume that we can always find the least favorable pair $(Q_{\infty},Q_{1})$. In Section VII, we provide several examples where this assumption is valid. Finally, if multiple pairs are satisfying (14), then we assume that any one pair has been chosen arbitrarily.
@@
$\min_{T}\;\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathcal{L}_{\texttt{{WADD}}}(T)$ $\geq\sup_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\min_{T}\;\mathcal{L}_{\texttt{{WADD}}}(T)$
$\geq\frac{\log\gamma}{\inf_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{\texttt{{KL}}}(P_{1}\|P_{\infty})}(1+o(1)),$ (15)
@@
$\frac{\log\gamma}{\inf_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{\texttt{{KL}}}(P_{1}\|P_{\infty})}$
provides a lower bound on the asymptotic performance of any algorithm. This lower bound can be achieved by a CUSUM-type algorithm $T^{*}$ if $\inf_{P_{\infty}\in\mathcal{G}_{\infty}}\mathbb{E}_{\infty}[T^{*}]\geq\gamma$ and the drift of this algorithm after the change is $\inf_{(P_{\infty},P_{1})\in\mathcal{G}_{\infty}\times\mathcal{G}_{1}}\mathbb{D}_{\texttt{{KL}}}(P_{1}\|P_{\infty})$. Since we cannot calculate the KL divergence for unnormalized and score-based models, we use the Fisher divergence as a proxy measure. We note also that these two divergences coincide for certain Gaussian families [15].
@@
$z(x)=\mathcal{S}_{\texttt{{H}}}(x,Q_{\infty})-\mathcal{S}_{\texttt{{H}}}(x,Q_{1}),$ (16)
@@
$Z(0)=0,\;\quad\quad Z(n)=(Z(n-1)+z(X_{n}))^{+},\;\forall n\geq 1.$
@@ VI Consistency and Delay and False Alarm Analysis of the RSCUSUM Algorithm
For the least favorable distributions $Q_{\infty},Q_{1}$ of Definition V.1, $\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})<\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1})$.
@@
If $\mathcal{G}_{\infty}$ contains only one element ($\mathcal{G}_{\infty}=\{P_{\infty}\}$), then clearly the least favorable pre-change distribution $Q_{\infty}$ is equal to $P_{\infty}$ and $\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})=0$. Then, Assumption VI.1 follows from the disjointness of $\mathcal{G}_{\infty},\mathcal{G}_{1}$ given in Assumption IV.3 and the regularity conditions given in Assumption IV.4.
@@
$\mathcal{G}_{\infty}=\biggl{\{}x\mapsto\sum_{i=1}^{m}\alpha_{i}p_{i}(x):\sum_{i=1}^{m}\alpha_{i}=1,\alpha_{i}\geq 0\biggr{\}}.$ (18)
If $\mathbb{D}_{\texttt{{F}}}(P_{i}\|Q_{\infty})<\mathbb{D}_{\texttt{{F}}}(P_{i}\|Q_{1})$ for all $1\leq i\leq m$, then Assumption VI.1 holds for any $P_{\infty}\in\mathcal{G}_{\infty}$.
@@
We use $C_{R}$ to denote the term $\mathbb{E}_{X\sim R}[\frac{1}{2}\left\|\nabla_{x}\log r(X)\right\|_{2}^{2}]$ for any distribution $R$. Then
$\begin{split}\mathbb{E}_{X\sim R}[z(X)]=\mathbb{E}_{X\sim R}[\mathcal{S}_{\texttt{{H}}}(X,Q_{\infty})-\mathcal{S}_{\texttt{{H}}}(X,Q_{1})]&=\mathbb{D}_{\texttt{{F}}}(R\|Q_{\infty})-C_{R}-\mathbb{D}_{\texttt{{F}}}(R\|Q_{1})+C_{R}\\ &=\mathbb{D}_{\texttt{{F}}}(R\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(R\|Q_{1}),\end{split}$
@@
$=\sum_{i=1}^{m}\alpha_{i}\bigg{(}\mathbb{D}_{\texttt{{F}}}(P_{i}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{i}\|Q_{1})\bigg{)}.$ (19)
It is given that $\mathbb{D}_{\texttt{{F}}}(P_{i}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{i}\|Q_{1})<0$ for each $1\leq i\leq m$ and that $\alpha_{i}\geq 0$ for all $1\leq i\leq m$. Thus, $\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1})$ can be written as the sum of $m$ nonpositive terms, at least one of which is strictly negative. ∎
@@ VI-A Reverse Triangle Inequality for Fisher Divergence
[…]
Clearly $\mathcal{L}(\xi)$ is minimized at $\xi=1$, and ${\partial\mathcal{L}(\xi)}/{\partial\xi}\mid_{\xi=1^{-}}\leq 0$. Letting $\mathcal{L}^{\prime}(\xi)={\partial\mathcal{L}(\xi)}/{\partial\xi}$, we have
$\mathcal{L}^{\prime}(\xi)=\int\frac{1}{2}\bigl{(}q_{1}-r_{1}\bigr{)}\big{\|}\nabla\log q_{\xi}-\nabla\log q_{\infty}\big{\|}^{2}dx\\ +\int q_{\xi}\nabla\left(\frac{q_{1}-r_{1}}{q_{\xi}}\right)^{T}\bigl{(}\nabla\log q_{\xi}-\nabla\log q_{\infty}\bigr{)}dx.$
@@
$\frac{1}{2}r_{1}\big{\|}\nabla\log q_{1}-\nabla\log q_{\infty}\big{\|}^{2}$ $=\frac{1}{2}r_{1}\big{\|}\nabla\log q_{1}-\nabla\log r_{1}\big{\|}^{2}+\frac{1}{2}r_{1}\big{\|}\nabla\log r_{1}-\nabla\log q_{\infty}\big{\|}^{2}$
$\quad+\underbrace{r_{1}\bigl{(}\nabla\log q_{1}-\nabla\log r_{1}\bigr{)}^{T}\bigl{(}\nabla\log r_{1}-\nabla\log q_{\infty}\bigr{)}}_{\text{term 1(a)}}.$ (21)
@@
$q_{1}\nabla\left(\frac{q_{1}-r_{1}}{q_{1}}\right)^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}\\ =r_{1}\bigl{(}\nabla\log q_{1}-\nabla\log r_{1}\bigr{)}^{T}\bigl{(}\nabla\log q_{1}-\nabla\log q_{\infty}\bigr{)}.$ (24)
@@
The results follows since ${\partial\mathcal{L}(\xi)}/{\partial\xi}\mid_{\xi=1^{-}}\leq 0$. ∎
@@ VI-B Consistency of the RSCUSUM Algorithm
$\mathbb{E}_{\infty}\left[z(X_{1})\right]$ $=\mathbb{E}_{\infty}\left[\mathcal{S}_{\texttt{{H}}}(X_{1},Q_{\infty})-\mathcal{S}_{\texttt{{H}}}(X_{1},Q_{1})\right]=\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1})<0,\;\text{and}$
@@
As in Theorem VI.3, we define
$C_{R}=\mathbb{E}_{X\sim R}\left[\frac{1}{2}\left\|\nabla_{x}\log r(X)\right\|_{2}^{2}\right]$
for any distribution $R$. Then
$\mathbb{E}_{\infty}[\mathcal{S}_{\texttt{{H}}}(X_{1},Q_{\infty})-\mathcal{S}_{\texttt{{H}}}(X_{1},Q_{1})]$ $=\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-C_{P_{\infty}}-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1})+C_{P_{\infty}}=\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{\infty}\|Q_{1}),$
@@
$\mathbb{E}_{1}[\mathcal{S}_{\texttt{{H}}}(X_{1},Q_{\infty})-\mathcal{S}_{\texttt{{H}}}(X_{1},Q_{1})]$ $=\mathbb{D}_{\texttt{{F}}}(P_{1}\|Q_{\infty})-C_{P_{1}}-\mathbb{D}_{\texttt{{F}}}(P_{1}\|Q_{1})+C_{P_{1}}\geq\mathbb{D}_{\texttt{{F}}}(Q_{1}\|Q_{\infty}),$
@@ VI-C False Alarm Analysis of the RSCUSUM Algorithm
$z_{\rho}(x)=\rho z(x)=\rho\bigr{(}\mathcal{S}_{\texttt{{H}}}(x,Q_{\infty})-\mathcal{S}_{\texttt{{H}}}(x,Q_{1})\bigr{)}.$ (27)
@@
for some $\rho$ which satisfies the inequality of Equation (26). To satisfy the constraint of $\mathbb{E}_{\infty}[T_{\texttt{{RSCUSUM}}}]\geq\gamma$, it is enough to set the threshold $\omega={(\log\gamma)}/{\rho}$.
@@
As a result, by setting $\omega={(\log\gamma)}/{\rho^{*}}$, we get
@@ VI-D Delay Analysis of the RSCUSUM Algorithm
$\mathcal{L}_{\texttt{{WADD}}}$ $(T_{\texttt{{RSCUSUM}}})\sim\mathcal{L}_{\texttt{{CADD}}}(T_{\texttt{{RSCUSUM}}})\sim\mathbb{E}_{1}[T_{\texttt{{RSCUSUM}}}]\sim\frac{\omega}{\mathbb{D}_{\texttt{{F}}}(P_{1}\|Q_{\infty})-\mathbb{D}_{\texttt{{F}}}(P_{1}\|Q_{1})}\lesssim\frac{\omega}{\mathbb{D}_{\texttt{{F}}}(Q_{1}\|Q_{\infty})},$
as $\omega\to\infty$.
Furthermore, if we let $\omega={(\log\gamma)}/{\rho}$ as in Theorem VI.7, then we have $\mathbb{E}_{\infty}[T_{\texttt{{RSCUSUM}}}]\geq\gamma$, and
$\mathcal{L}_{\texttt{{WADD}}}(T_{\texttt{{RSCUSUM}}})\sim\mathcal{L}_{\texttt{{CADD}}}(T_{\texttt{{RSCUSUM}}})\lesssim\frac{\log\gamma}{\rho\mathbb{D}_{\texttt{{F}}}(Q_{1}\|Q_{\infty})},$
as $\gamma\to\infty$.
@@
In the above theorem, we have used the notation $g(c)\lesssim h(c)$ as $c\to c_{0}$ to indicate that $\lim\sup{g(c)}/{h(c)}\leq 1$ as $c\to c_{0}$ for any two functions $c\mapsto g(c)$ and $c\mapsto h(c)$.
Theorem VI.8 implies that the expected detection delay (EDD), $\mathbb{E}_{1}[T_{\texttt{{RSCUSUM}}}]$, increases at most linearly as the stop threshold $\omega$ increases for large values of $\omega$. Again, since the bound is valid for all those $\rho$ values for which $\mathbb{E}_{\infty}[\exp(z_{\rho}(X_{1}))]\leq 1$, the bound is valid for the biggest such number, which is $\rho^{*}$. This gives
$\mathcal{L}_{\texttt{{WADD}}}(T_{\texttt{{RSCUSUM}}})\sim\mathcal{L}_{\texttt{{CADD}}}(T_{\texttt{{RSCUSUM}}})\lesssim\frac{\log\gamma}{\rho^{*}\mathbb{D}_{\texttt{{F}}}(Q_{1}\|Q_{\infty})},\;\;$ (30)
as $\gamma\to\infty$.